OpenStudy (anonymous):

If the cos 30° = square root of 3/2 , then the sin 60° = _____.

OpenStudy (anonymous):

\[\sin \theta=\cos \left( 90-\theta \right) \] \[put \theta=60 \]

OpenStudy (anonymous):

so its 0 ?

OpenStudy (anonymous):

How ?

OpenStudy (campbell_st):

here is the triangle  what do you think sin(60) is ?